OpenStudy (anonymous):

integral of x/(x^4 + 25)

OpenStudy (turingtest):

\[u=x^2+25\implies du=?\]

OpenStudy (anonymous):

2x dx, so dx = 1/2 du

OpenStudy (anonymous):

1/2 integral of du/u

OpenStudy (turingtest):

xdx=1/2du

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

wait I typed it wrong sorry, it's x^4

OpenStudy (turingtest):

\[\int\frac{xdx}{x^4+25}\]?

OpenStudy (anonymous):

yes, do I somehow use the derivative of arctan?

OpenStudy (turingtest):

yeah that should work

OpenStudy (turingtest):

u=x^2

OpenStudy (anonymous):

but derivative of arctan is 1/(x^2 + 1), so how do I get rid of the 25?

OpenStudy (turingtest):

\[\int\frac{dx}{x^2+a^2}=\frac1a\tan^{-1}(\frac xa)+C\]if we want to prove that it will take a minute

OpenStudy (anonymous):

1/5 arctan (x^2/5)?

OpenStudy (turingtest):

careful, you have to make the u-sub first

OpenStudy (turingtest):

very close

OpenStudy (anonymous):

1/10 not 1/5 since 1/2 du = dx?

OpenStudy (turingtest):

exactly, and don't forget +C if it's indefinite

OpenStudy (anonymous):